1. A 50 kg sled sits motionless on a ramp with a slope of 15o. What is the force of friction on the sled? 

(a) 47 N 

(b) 127 N 

(c) 155 N 

(d) 490 N 



1
Expert's answer
2020-11-05T07:34:18-0500
"F=mg\sin{15}\\F=(50)(9.8)\sin{15}=127\ N"

Answer: (b) 127 N